# Distillation-Free One-Step Diffusion for Real-World Image Super-Resolution

### v2

@@ 3.2 Noise-Aware Discriminator (NAD)
where $\hat{z}_{H}$ is computed as: $\hat{z}_{H}=\frac{z_{L}-\sqrt{1-\bar{\alpha}_{T}}\,\epsilon_{\theta}(z_{L};T)}{\sqrt{\bar{\alpha}_{T}}}$, and $F(\cdot,t)$ denotes the forward diffusion process of $\cdot$ at time step $t\in[0,T]$, specifically,
$F(z,t)=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\text{with}~{}\epsilon{\sim}\mathcal{N}(0,\mathbf{I}).$ (4)
@@ 3.3 Edge-Aware DISTS
To further enhance the quality of the generated images, we aim to incorporate perceptual loss. Most image reconstruction methods utilize LPIPS (Learned Perceptual Image Patch Similarity) (Zhang et al., 2018a) as the perceptual loss. However, to better preserve image texture details and alleviate pseudo-textures in the reconstruction under higher noise levels, we need to focus on the textures on HR images. DISTS (Deep Image Structure and Texture Similarity) (Ding et al., 2020) can compute the structural and textural similarity of images, aligning with human subjective perception of image quality. Furthermore, regions with rich textures or details often exhibit strong edge information. Leveraging image edge information effectively enhances texture quality. Based on this, we propose a novel perceptual loss, termed Edge-Aware DISTS (EA-DISTS). This perceptual loss simultaneously evaluates the structure and texture similarity of the reconstructed and HR images and their edges, thereby enhancing texture detail restoration.
@@
$\mathcal{L}_{\text{EA-DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H})=\mathcal{L}_{\text{DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H})+\mathcal{L}_{\text{DISTS}}(\mathcal{S}(\mathcal{G}_{\theta}(x_{L})),\mathcal{S}(x_{H})),$ (5)
@@
where * denotes the convolution operation. This computation results in an edge map that highlights the structural and textural details of the image.
@@
3.4 Distillation-Free Training
Here, we summarize the whole distillation-free one-step diffusion model training process. As described in Section 3.1, within the generator component, DFOSD obtains $\hat{z}_{H}$ and the decoded high-resolution image $\hat{x}_{H}$ through one-step sampling. The generator then updates its parameters by computing the spatial loss $\mathcal{L}_{\text{spatial}}$ in pixel space between the generated image and the ground truth, as well as the adversarial loss $\mathcal{L}_{\mathcal{G}}$ derived from the discriminator in the latent space (Eq. 2). The loss function for updating the generator is defined as $\mathcal{L}_{\text{spatial}}+\lambda_{1}\mathcal{L}_{\mathcal{G}}$. Specifically, we employ a weighted sum of Mean Squared Error (MSE) loss and perceptual loss to define the spatial loss:
$\mathcal{L}_{\text{spatial}}(\mathcal{G}_{\theta}(x_{L}),x_{H})=\mathcal{L}_{\text{MSE}}(\mathcal{G}_{\theta}(x_{L}),x_{H})+\lambda_{2}\mathcal{L}_{\text{EA-DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H}),$ (8)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters used to balance the contributions of each loss component.
For discriminator training, we utilize paired training features, where each pair consists of a negative sample feature $\hat{z}_{H}$ (generated by the generator) and the corresponding real image’s latent representation $z_{H}$ as a positive one. Using Eq. 3, we compute the adversarial loss $\mathcal{L}_{\mathcal{D}}$ to update the discriminator’s parameters. Furthermore, the discriminator can be initialized with weights from more powerful pre-trained models, such as SDXL (Podell et al., 2023), to achieve superior performance.
This distillation-free training approach allows our DFOSD to overcome the limitations imposed by multi-step diffusion models, enhancing generator performance without increasing its parameter count or compromising efficiency. Additionally, the integration of a robust discriminator initialized with advanced pre-trained models ensures that the generator receives high-quality feedback, facilitating the production of more realistic and detailed high-resolution images.
Multi-step Diffusion One-step Diffusion
Datasets Metrics StableSR-s200 DiffBIR-s50 SeeSR-s50 ResShift-s15 SinSR-s1 OSEDiff-s1 DFOSD-s1
NIQE↓ 4.8927 3.9472 4.5403 7.3495 5.7467 4.3443 3.9255
MUSIQ↑ 60.53 68.02 66.37 56.18 61.62 67.31 69.21
ManIQA↑ 0.5570 0.6309 0.6118 0.5004 0.5362 0.6148 0.6402
RealSR ClipIQA↑ 0.5140 0.7295 0.6822 0.5848 0.6927 0.6827 0.6683
NIQE↓ 4.9852 4.1218 4.6891 6.7303 5.6642 4.2245 3.9580
MUSIQ↑ 58.89 71.23 69.79 59.36 64.22 69.04 69.69
ManIQA↑ 0.5269 0.6371 0.6018 0.5071 0.5338 0.6024 0.6215
RealSet65 ClipIQA↑ 0.5609 0.7734 0.7004 0.6331 0.7263 0.6874 0.6843
NIQE↓ 5.3139 3.0885 4.1390 7.0159 5.5639 4.3661 4.1682
MUSIQ↑ 34.68 36.18 34.51 30.52 32.79 37.22 40.30
ManIQA↑ 0.4675 0.5985 0.5758 0.4210 0.4755 0.5797 0.5703
DRealSR ClipIQA↑ 0.5208 0.7568 0.6746 0.5884 0.7231 0.7540 0.6914
Table 1: Quantitative no-reference (NR) metrics comparison with state-of-the-art DM-based methods for Real-ISR ($\times$4). The best and second-best results of each metric within both multi-step and one-step diffusion-based methods are highlighted in red and blue, respectively.
@@ 4 Experiments
We conduct comprehensive experiments to validate the effectiveness of DFOSD in real-world image super-resolution (Real-ISR). We provide a detailed introduction of our experimental setup in Section 4.1. In Section 4.2, we evaluate our method on three challenging real-world datasets: RealSR (Cai et al., 2019), RealSet65 (Yue et al., 2024), and DRealSR (Wei et al., 2020), and compare it against the current state-of-the-art methods. In Section 4.3, we carry out comprehensive ablation studies to validate the effectiveness and robustness of our proposed approach.
@@ 4.1 Experimental Settings
Datasets. We train DFOSD on a self-collected dataset comprising 200K high-quality images. During training, we randomly crop patches of size 512$\times$512 pixels from these images. To generate low-resolution (LR) and high-resolution (HR) pairs for training, we apply the Real-ESRGAN degradation pipeline. We conduct extensive evaluations of DFOSD on multiple real-world datasets, including RealSR (Cai et al., 2019), RealSet65 (Yue et al., 2024), and DRealSR (Wei et al., 2020). To avoid potential biases and ensure a fair comparison, we evaluate our model and all other methods by using the whole images from each dataset. We assess image quality without any cropping (e.g., random crop, central crop) that might make the evaluation results randomly and hard to reproduce.
Non-Diffusion Multi-step Diffusion One-step Diffusion
Datasets Metrics Real-ESRGAN SwinIR StableSR-s50 DiffBIR-s50 SeeSR-s50 ResShift-s15 SinSR-s1 OSEDiff-s1 DFOSD-s1
PSNR↑ 30.55 28.31 30.31 25.91 28.35 26.42 27.33 24.20 26.47
SSIM↑ 0.8571 0.8273 0.8394 0.6190 0.8052 0.7310 0.7237 0.7355 0.7838
LPIPS↓ 0.3843 0.2736 0.2818 0.5347 0.3031 0.4582 0.4444 0.3429 0.3149
DRealSR DISTS↓ 0.2034 0.1387 0.1428 0.2387 0.1665 0.2382 0.2262 0.1763 0.1547
PSNR↑ 27.57 27.34 26.28 24.87 26.20 25.45 25.83 24.57 24.60
SSIM↑ 0.7741 0.7862 0.7733 0.6486 0.7555 0.7246 0.7183 0.7202 0.7221
LPIPS↓ 0.2729 0.2515 0.2622 0.3834 0.2806 0.3727 0.3641 0.3036 0.3031
RealSR DISTS↓ 0.1542 0.1583 0.2147 0.2015 0.1784 0.2344 0.2193 0.1808 0.1775
Table 2: Quantitative FR metrics comparison for Real-ISR ($\times$4). The best and second-best results within both multi-step and one-step diffusion-based methods are highlighted in red, blue, respectively.
Refer to caption Canon_001 PSNR / SSIM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic BSRGAN Real-ESRGAN RealSR_JPEG SwinIR DFOSD (ours) 29.21 / 0.8687 29.15 / 0.8940 28.16 / 0.8862 29.58 / 0.8635 28.85 / 0.9028 26.48 / 0.8254
Refer to caption Canon_018 PSNR / SSIM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic BSRGAN Real-ESRGAN RealSR_JPEG SwinIR DFOSD (ours) 29.43 / 0.8456 28.31 / 0.8389 26.68 / 0.8318 29.31 / 0.8321 27.72 / 0.8428 24.46 / 0.7749
Figure 5: Visual comparison ($\times$4) of DFOSD with GAN-based and Transformer-based methods. Cannon_001 contains the letters ‘eḋu’. Cannon_018 contains the structures of tower windows. Although GAN-based approaches achieve higher PSNR and SSIM scores, their generated images exhibit less realistic and detailed textures compared to DFOSD. Those quantitative and visual comparisons indicate that higher PSNR and SSIM values do not mean better visual quality.
Implementation Details. We adopt Stable Diffusion (SD) 2.1-base as the backbone for training DFOSD, setting both the rank and scaling factor $\alpha$ of LoRA to 16 in the generator and discriminator. The model is trained using the AdamW optimizer with learning rates of 5$\times$$10^{-5}$ for both generator and discriminator. We utilize a learnable text embedding as the conditional input for the SD UNet, without any prompts, and remove the text encoder. Training is performed with a batch size of 16 over 100K iterations with 4 NVIDIA A100-40GB GPUs.
@@
Compared Methods. We compare our DFOSD with state-of-the-art diffusion model (DM)-based methods for real image super-resolution (Real-ISR), as well as other prominent approaches, including GAN-based and Transformer-based methods. The DM-based methods encompass multi-step diffusion models, such as StableSR (Wang et al., 2024a), ResShift (Yue et al., 2024), DiffBIR (Lin et al., 2024), and SeeSR (Wu et al., 2024b), alongside recently proposed one-step diffusion models like SinSR (Wang et al., 2024b) and OSEDiff (Wu et al., 2024a). OSEDiff is the current top-performing one-step diffusion Real-ISR method. Other methods include GAN-based approaches, such as BSRGAN (Zhang et al., 2021), RealSR-JPEG (Ji et al., 2020), and Real-ESRGAN (Wang et al., 2021b), as well as Transformer-based method SwinIR (Liang et al., 2021).
Evaluation Metrics. To comprehensively assess the performance of each method, we employ four full-reference (FR) and four no-reference (NR) image quality metrics. The FR metrics consists of Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018a), and Deep Image Structure and Texture Similarity (DISTS) (Ding et al., 2020). PSNR measures pixel-wise differences, while SSIM evaluates structural similarity. Both PSNR and SSIM are computed on the Y channel in the YCbCr color space. LPIPS assesses perceptual similarity using deep neural network features. DISTS combines structural and textural comparisons. The NR metrics include Naturalness Image Quality Evaluator (NIQE) (Zhang et al., 2015), Multi-scale Image Quality Transformer (MUSIQ) (Ke et al., 2021), Multi-scale Attention-based Image Quality Assessment (ManIQA) (Yang et al., 2022), and ClipIQA (Wang et al., 2023a). NIQE evaluates image quality based on statistical features. MUSIQ captures multi-scale distortions using Transformers. ManIQA employs attention mechanisms to assess quality. ClipIQA leverages pre-trained models like CLIP to align quality assessments with human perception.
@@ 4.2 Comparison with State-of-the-Art Methods
Quantitative Results. Tables 1 and 2 provide quantitative comparisons of the methods across the three datasets. DFOSD achieves either the best or second-best performance on the majority of metrics across all datasets when compared with other one-step diffusion methods. Although GAN-based methods outperform diffusion-based methods in terms of PSNR and SSIM, they generally exhibit poorer performance on NR metrics. Detailed comparisons of NR metrics and visual results for non-diffusion-based methods are provided in the supplementary material. Despite the higher FR metrics achieved by GAN-based and Transformer-based methods, their visual results are significantly inferior to those of DFOSD. Figure 5 illustrates several examples, further highlighting the limitations of full-reference metrics in accurately evaluating image quality. This underscores the necessity for more effective approaches to assess the quality of generated images.
StableSR DiffBIR SeeSR ResShift SinSR OSEDiff DFOSD (ours)
# Step 200 50 50 15 1 1 1
Inference Time / s 11.50 7.79 5.93 0.71 0.16 0.35 0.11
# Total Param / M 1.4$\times$$10^{3}$ 1.6$\times$$10^{3}$ 2.0$\times$$10^{3}$ 173.8 173.8 1.4$\times$$10^{3}$ 966.3
# MACs / G 75,812 24,528 32,336 4,903 2,059 2,269 2,132
@@
Dataset NIQE↓ MUSIQ↑ ManIQA↑ ClipIQA↑
LSDIR + 10K FFHQ 3.9264 67.26 0.6140 0.6397
Our Dataset 3.9255 69.21 0.6402 0.6683
Table 4: Quantitative comparison ($\times$4) on RealSR. Our DFOSD is trained on different datasets.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff DFOSD (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff DFOSD (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff DFOSD (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff DFOSD (ours)
Figure 6: Visual comparisons ($\times$4) on Real-ISR task.
Visual Results. Figure 6 presents a visual comparison of various diffusion-based Real-ISR methods. As observed, most existing methods struggle to generate realistic details and often produce incorrect content in certain regions of the image due to noise artifacts. Notably, our DFOSD demonstrates a significant advantage over others, particularly in the restoration of textual content. Additional visual comparison results are provided in the supplementary material.
Complexity Analysis. Table 3 presents a complexity comparison of DM-based Real-ISR methods, including the number of inference steps, inference time, parameter numbers, and MACs (Multiply-Accumulate Operations). All methods are evaluated on an NVIDIA A100 GPU. DFOSD achieves the fastest inference speed among all DM-based methods. Furthermore, since we do not employ a text encoder or other additional modules (such as DAPE used by OSEDiff and SeeSR, and ControlNet used by DiffBIR), our DFOSD has the smallest number of model parameters during inference among Stable Diffusion (SD)-based methods, reducing the parameters by $33\%$ compared to OSEDiff.

### v3

@@ 3.2 Distillation with Large Diffusion Discriminator
where $F(\cdot,t)$ denotes the forward diffusion process of $\cdot$ at time step $t\in[0,T]$, specifically,
$F(z,t)=\sqrt{\bar{\alpha}_{t}}\,z+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\text{with}\leavevmode\nobreak\ \epsilon{\sim}\mathcal{N}(0,\mathbf{I}).$ (4)
Relation to Diffusion GAN Methods. Recently, many methods have combined GANs and Diffusion models, achieving success in image generation and other fields. Diffusion-GAN [48] uses a timestep-dependent discriminator to guide the generator’s training. DDGAN [53] employs a multimodal conditional GAN to achieve large-step denoising. ADD [39] and LADD [38] both introduce discriminators to enhance the generative quality of one-step student diffusion models. These discriminators operate in pixel space and latent space, respectively. Both our method and existing Diffusion GAN methods demonstrate the potential of adversarial training in diffusion models. However, there are several essential differences. First, we use a pre-trained multi-step diffusion model as the discriminator, leveraging the prior of large-scale models to guide the training process. Second, we explore the impact of scaling. Figure 3 shows the performance of different SD models as discriminators. It reveals that large-scale diffusion discriminators break through the performance limits of pre-trained diffusion models, achieving superior results.
@@ 3.3 Edge-Aware DISTS
To further enhance the quality of the generated images, we aim to incorporate perceptual loss. Most image reconstruction methods utilize LPIPS [65] as the perceptual loss. However, to better preserve image texture details and alleviate pseudo-textures in the reconstruction under higher noise levels, we need to focus on the textures on HR images. DISTS [7] can compute the structural and textural similarity of images, aligning with human subjective perception of image quality. Furthermore, regions with rich textures or details often exhibit strong edge information. Leveraging image edge information effectively enhances texture quality. Based on this, we propose a novel perceptual loss, termed Edge-Aware DISTS (EA-DISTS). This perceptual loss simultaneously evaluates the structure and texture similarity of the reconstructed and HR images and their edges, thereby enhancing texture detail restoration.
@@
$\mathcal{L}_{\text{EA-DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H})=$
$\mathcal{L}_{\text{DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H})+\mathcal{L}_{\text{DISTS}}(\mathcal{S}(\mathcal{G}_{\theta}(x_{L})),\mathcal{S}(x_{H})),$ (5)
@@
where * denotes the convolution operation.
@@
3.4 Overall Training Scheme
Here, we summarize the whole one-step diffusion model training process. As described in Sec. 3.1, within the generator component, D3SR obtains $\hat{z}_{H}$ and the decoded high-resolution image $\hat{x}_{H}$ through one-step sampling. The generator then updates its parameters by computing the spatial loss $\mathcal{L}_{\text{spatial}}$ in pixel space between the generated image and the ground truth, as well as the adversarial loss $\mathcal{L}_{\mathcal{G}}$ derived from the discriminator in the latent space (Eq. 2). The loss function for updating the generator is defined as $\mathcal{L}_{\text{spatial}}+\lambda_{1}\mathcal{L}_{\mathcal{G}}$. Specifically, we employ a weighted sum of Mean Squared Error (MSE) loss and perceptual loss to define the spatial loss:
$\mathcal{L}_{\text{spatial}}(\mathcal{G}_{\theta}(x_{L}),x_{H})=$
$\mathcal{L}_{\text{MSE}}(\mathcal{G}_{\theta}(x_{L}),x_{H})+\lambda_{2}\mathcal{L}_{\text{EA-DISTS}}(\mathcal{G}_{\theta}(x_{L}),x_{H}),$ (8)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters used to balance the contributions of each loss component. The supplementary material provide a detailed description of the algorithm’s pseudocode and the values of the hyperparameters.
For discriminator training, we utilize paired training features, where each pair consists of a negative sample feature $\hat{z}_{H}$ and the corresponding real image’s latent representation $z_{H}$ as a positive one. Using Eq. 3, we compute the adversarial loss $\mathcal{L}_{\mathcal{D}}$ to update the discriminator’s parameters. Furthermore, the discriminator can be initialized with weights from more powerful pre-trained models, such as SDXL [35], to achieve superior performance.
This training approach allows our D3SR to overcome the limitations imposed by teacher models, enhancing generator performance without increasing its parameter count or compromising efficiency. Additionally, the integration of a robust discriminator initialized with advanced pre-trained models ensures that the generator receives high-quality feedback, facilitating the production of more realistic and detailed high-resolution images.
@@ 4 Experiments
We conduct comprehensive experiments to validate the effectiveness of D3SR in real-world image super-resolution (Real-ISR). We provide a detailed introduction of our experimental setup in Sec. 4.1. In Sec. 4.2, we evaluate our method and compare it against the current state-of-the-art methods. In Sec. 4.3, we carry out comprehensive ablation studies to validate the effectiveness and robustness of our proposed approach.
Table 1: Quantitative results ($\times$4) on the Real-ISR testset with ground truth. The best and second-best results are colored red and blue.
Dataset Method PSNR↑ SSIM↑ LPIPS↓ DISTS↓ NIQE↓ MUSIQ↑ MANIQA↑
SinSR [45] 25.83 0.7183 0.3641 0.2193 5.746 61.62 0.5362
OSEDiff [51] 24.57 0.7202 0.3036 0.1808 4.344 67.31 0.6148
RealSR ADDSR 25.23 0.7295 0.2990 0.1852 5.223 63.08 0.5457
[…]
D3SR 23.98 0.7133 0.3150 0.1906 3.747 68.69 0.6436
SinSR [45] 22.55 0.5405 0.4390 0.2033 5.620 62.25 0.5011
OSEDiff [51] 23.10 0.6127 0.3447 0.1750 3.583 66.62 0.5530
DIV2K-val ADDSR 22.74 0.6007 0.3961 0.1974 4.270 62.08 0.5118
[…]
D3SR 21.46 0.5877 0.3873 0.1663 3.215 69.71 0.5848
Table 2: Quantitative results ($\times$4) on RealSet65 testset. The best and second-best results are colored red and blue. In the one-step diffusion models, the best metric is bolded.
Method NIQE↓ MUSIQ↑ MANIQA↑
StableSR-s200 4.985 58.89 0.5269
DiffBIR-s50 4.122 71.23 0.6371
SeeSR-s50 4.689 69.79 0.6018
ResShift-s15 6.730 59.36 0.5071
ADDSR-s4 5.390 68.97 0.6075
[…]
OSEDiff-s1 4.224 69.04 0.6024
ADDSR-s1 5.207 64.22 0.5258
[…]
D3SR-s1 4.244 71.96 0.6315
@@ 4.1 Experimental Settings
Datasets. We train D3SR on both public datasets and our collected large-scale high-quality dataset. The public datasets include LSDIR [22] and the first 10k images from FFHQ [18], totaling 95k images. Our collected dataset contains 200k high-quality images. During training, we randomly crop patches of size 512$\times$512 pixels from these images. To get low-resolution (LR) and high-resolution (HR) pairs for training, we apply the Real-ESRGAN [44] degradation pipeline. We conduct extensive evaluations of D3SR on a synthetic dataset DIV2K-val [1] and two real-world datasets, including RealSR [3] and RealSet65 [61]. In DIV2K-val, we use the Real-ESRGAN degradation pipeline to synthesize the corresponding LR images. We evaluate our model and all other methods by using the whole images from each dataset.
@@
Compared Methods. We compare our D3SR with state-of-the-art DM-based methods for real image super-resolution (Real-ISR), as well as other prominent approaches, including GAN-based and Transformer-based methods. The DM-based methods include multi-step diffusion models, such as StableSR [42], ResShift [61], DiffBIR [25], and SeeSR [52], alongside recently proposed one-step diffusion models like SinSR [45], OSEDiff [51], and AddSR [54]. Other methods include GAN-based approaches, such as BSRGAN [62], RealSR-JPEG [15], Real-ESRGAN [44], LDL [24], and FeMASR [4], as well as Transformer-based method SwinIR [23].
Evaluation Metrics. To comprehensively assess the performance of each method, we employ four full-reference (FR) and three no-reference (NR) image quality metrics. The FR metrics consists of PSNR, SSIM, LPIPS [65], and DISTS [7]. Both PSNR and SSIM are computed on the Y channel in the YCbCr color space. The NR metrics include NIQE [63], MUSIQ [19], and ManIQA [55].
Implementation Details. We initialize the generator network with the SD 2.1-base parameters and the discriminator network with partial parameters from SDXL. We set both the rank and scaling factor $\alpha$ of LoRA to 16 in the generator and discriminator. We use the AdamW optimizer and set the learning rate for both the generator and discriminator to 5e-5. Training is performed with a batch size of 8 over 100K iterations with 4 NVIDIA A100-40GB GPUs.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff D3SR (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff D3SR (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Bicubic Real-ESRGAN SwinIR StableSR DiffBIR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ResShift SeeSR SinSR OSEDiff D3SR (ours)
Figure 5: Visual comparisons ($\times$4) on Real-ISR task (RealSR [3] dataset).
@@ 4.2 Comparison with State-of-the-Art Methods
Table 3: Quantitative results ($\times$4) on the Real-ISR testset with ground truth. The best and second-best results are colored red and blue. Trained on public datasets.
Dataset Method PSNR↑ SSIM↑ LPIPS↓ DISTS↓ NIQE↓ MUSIQ↑ MANIQA↑
StableSR-s200 26.28 0.7733 0.2622 0.1583 4.892 60.53 0.5570
DiffBIR-s50 24.87 0.6486 0.3834 0.2015 3.947 68.02 0.6309
SeeSR-s50 26.20 0.7555 0.2806 0.1784 4.540 66.37 0.6118
RealSR ResShift-s15 25.45 0.7246 0.3727 0.2344 7.349 56.18 0.5004
ADDSR-s4 23.15 0.6662 0.3769 0.2353 5.256 66.54 0.6581
[…]
D3SR-s1 23.98 0.7133 0.3150 0.1906 3.747 68.69 0.6436
StableSR-s200 23.68 0.6270 0.4167 0.2023 4.602 49.51 0.4774
DiffBIR-s50 22.33 0.5133 0.4681 0.1889 3.156 70.07 0.6307
SeeSR-s50 23.21 0.6114 0.3477 0.1706 3.591 67.99 0.5959
DIV2K-val ResShift-s15 23.55 0.6023 0.4088 0.2228 6.870 56.07 0.4791
ADDSR-s4 22.08 0.5578 0.4169 0.2145 4.738 68.26 0.5998
[…]
D3SR-s1 21.46 0.5877 0.3873 0.1663 3.215 69.71 0.5848
Quantitative Results. Tables 1, 2 and 3 provide quantitative comparisons of the methods across the three datasets. D3SR achieves the best performance among all no-reference (NR) metrics in one-step diffusion methods. Recently, some studies [60] have pointed out that reference-based metrics such as PSNR and SSIM cannot accurately reflect the performance of diffusion-based Real-ISR methods. We discuss this in the supplementary materials. We compare the quantitative and qualitative results, including those of GAN-based methods. We find that, compared to diffusion-based methods, GAN-based methods generally achieve higher PSNR and SSIM. However, the image quality of these methods is not as high.
Visual Results. Figure 5 presents a visual comparison of various diffusion-based Real-ISR methods. As observed, most existing methods struggle to generate realistic details and often produce incorrect content in certain regions of the image due to noise artifacts. Notably, our D3SR demonstrates a significant advantage over others, particularly in the restoration of textual content. Additional visual comparison results are provided in the supplementary material.
Complexity Analysis. Table 4 presents a complexity comparison of Stable Diffusion (SD)-based Real-ISR methods, including the number of inference steps, inference time, parameter numbers, and MACs (Multiply-Accumulate Operations). All methods are evaluated on an NVIDIA A100 GPU. D3SR achieves the fastest inference speed among all SD-based methods. Furthermore, our method supports using a fixed text embedding as the generation condition. Therefore, we do not require CLIP and other additional modules (such as DAPE used by OSEDiff and SeeSR, and ControlNet used by DiffBIR) for inference. Our D3SR has the smallest number of model parameters during inference among Stable Diffusion (SD)-based methods, reducing the parameters by $33\%$ compared to OSEDiff.
@@
StableSR DiffBIR SeeSR OSEDiff D3SR
# Step 200 50 50 1 1
Inference Time / s 11.50 7.79 5.93 0.35 0.11
# Total Param / M 1.4$\times$$10^{3}$ 1.6$\times$$10^{3}$ 2.0$\times$$10^{3}$ 1.4$\times$$10^{3}$ 966.3
# MACs / G 75,812 24,528 32,336 2,269 2,132